OpenStudy (anonymous):

The odds against Lounging Larry winning the horse race are 3:8. Determine the probability that Lounging Larry wins the horse race. 3/8 8/3 3/11 8/11 5/11

OpenStudy (anonymous):

\[\frac{8}{11}\]

OpenStudy (yaoliang):

If the odds of Lounging Larry winning the race are 3:8, then his chances are 3 out of 3+8, or 3 out of 11. Thus, his probability is 3/11..

OpenStudy (anonymous):

Thanks for the help

OpenStudy (yaoliang):

welcome

OpenStudy (anonymous):

careful question is worded funny 3:8 is odd AGAINST winning, so probability of not winning is \[\frac{3}{11}\] and so probabilituy of winning is \[\frac{8}{11}\]